OpenStudy (anonymous):

12-14. The table shows the percentage of total athletes at a school that play certain sports. Each athlete ONLY plays one sport. TWENTY athletes play baseball. How many athletes are in the school? Sport - Percentage: Baseball - ?% (unknown) Football - 30% Other - 45% 12. What percentage of athletes play baseball? 13. How many total athletes are there at the school? 14. How many athletes play football? Thank u so much! My older sister found a lot of help on here so i'm hoping for the same thing too! u all are great, thank u for answering, u are truly helping me!

OpenStudy (kropot72):

The percentage that play baseball = 100 - (30 + 45) = ?%

OpenStudy (anonymous):

@kropot72 Is it 25?

OpenStudy (kropot72):

Good work! You are right. You are given that 20 athletes play baseball. We now know that 25% of the total number of athletes at the school play baseball. Therefore 20 athletes are 25% of the total number of athletes. To get the total number of athletes at the school: Divide 20 by 25 and multiply the result by 100: \[Total\ number\ of\ athletes=\frac{20}{25}\times 100=you\ can\ calculate\]

OpenStudy (anonymous):

@kropot72 80? :)

OpenStudy (kropot72):

Good! There is a total of 80 athletes at the school. Now to answer the last part you need to find 30% of 80.

OpenStudy (anonymous):

@kropot72 How do I do that again? :)

OpenStudy (kropot72):

\[Number\ of \ athletes\ that\ play\ football=80\times \frac{30}{100}=?\]

OpenStudy (anonymous):

@kropot72 24? :D

OpenStudy (kropot72):

Correct again! Good work :)

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (kropot72):

You're welcome :)